OpenStudy (anonymous):

help on one 8th grade question? I really need help

OpenStudy (anonymous):

@Vocaloid if your on...can you help me with just one question?

OpenStudy (anonymous):

A container holds a combination of dimes and quarters. There are 178 coins in all, and their total value is $29.80. The system of equations models this situation, where x is the number of dimes and y is the number of quarters. x+y=178 0.1x+0.25y=29.8 How many dimes and quarters are in the container? A. 123 dimes and 55 quarters B. 114 dimes and 64 quarters C. 102 dimes and 76 quarters D. 98 dimes and 80 quarters

OpenStudy (anonymous):

hold on im solving

OpenStudy (anonymous):

ok :D

OpenStudy (anonymous):

sorry for taking so long :(

OpenStudy (anonymous):

its fine :)

OpenStudy (anonymous):

got it its D

OpenStudy (anonymous):

wait...how did you find out?

OpenStudy (anonymous):

x + 80 = 178 x = 178 - 80 = # of quarters

OpenStudy (anonymous):

oh...thanks :)

OpenStudy (anonymous):

happy to help

OpenStudy (anonymous):

crap...it was wrong :( I got a 60%

OpenStudy (anonymous):

I need to take it again...now I "really really" need to pass this... :(